OpenStudy (anonymous):

When a battery is connected to a 2.0 microfarad capacitor, it places an equal but opposite charge on each plate, of magnitude 4.0×10−5 coulombs. How much would the same battery charge a 3.6e−6 farad capacitor?

OpenStudy (anonymous):

3.6e-6 farad(let it be C) using V=Q/c find V of battery= say, V1.so simply use Q=CV again and find Q usinng the fact that C is our assumed value...does that help???